Mr. Granger is 10 years more than twice as old as his son. However, last year he was 4 years less than 3 times as old as his son. How old are Mr. Granger and his son?
Answer by fractalier(2101)

Let x be the son's age NOW (always set up variables in the NOW).
Then Granger's age NOW is 10 + 2x.
Last year, he was 9 + 2x. The son was x - 1. So
(9 + 2x) + 4 = 3(x - 1)
2x + 13 = 3x - 3
x = 16
Granger = 10 + 2x = 10 + 32 = 42